OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1 = -12 and a27 = 66

OpenStudy (anonymous):

i keep coming up with 126 but my options are 70,72,111,114. PLEASE HELP

OpenStudy (anonymous):



OpenStudy (jhannybean):

\[\large a_{n}= a_{1}+(n-1)d\]

OpenStudy (anonymous):

r=3

OpenStudy (anonymous):

111

OpenStudy (anonymous):

find r and then just plug in that first equation

OpenStudy (jhannybean):

\[\large 66 =-12+(27-1)d \]\[\large 66= -12 +26d\]\[\large 78 = 26d\]\[\large d= 3\] \[\large a_{42} = -12 + 3(42-1)\]\[\large a_{42} = -12 + 123\]\[\large a_{42} = 111\]

OpenStudy (anonymous):

i have no idea what i was doing wrong... but thank you